OpenStudy (anonymous):

says find the least common denominator of each group of fractions .some groups will include more than two fractions.they are 1/2and 1/3,and 1/5?besides 30

OpenStudy (anonymous):

Since all of those denominators are prime, the LCD of all of them will be the product of all of them. It must be 30.

OpenStudy (diyadiya):

30

OpenStudy (anonymous):

wat will the fractions look like

OpenStudy (anonymous):

You have to multiply each fraction by 1= a/a where a is the factors it's missing.

OpenStudy (anonymous):

i understand for each the denominator is 30 so ,wat are the other fraction numerators ?

OpenStudy (anonymous):

Well what version of 1 = a/a do you need to multiply 1/2 by to make the denominator 30?

OpenStudy (anonymous):

\[\frac{1}{2} \cdot \frac{a}{a} = \frac{a}{30}\] What must a be?

OpenStudy (anonymous):

\begin{align}\frac{1}{3}\cdot \frac{b}{b} &= \frac{b}{30}\\ \frac{1}{5} \cdot \frac{c}{c} &= \frac{c}{30}\end{align} What must b and c be?

OpenStudy (diyadiya):

1/2=15/30 1/3=10/30 1/5=6/30